# A module-theoretic interpretation of quantum expansion formula

### v1

@@ Abstract.
We provide a module-theoretic interpretation of the expansion formula given in [11], which defines a map on perfect matchings to compute the expansion of quantum cluster variables in quantum cluster algebras arising from unpunctured surfaces. In addition, we present a multiplication formula for string modules with one-dimensional extension spaces, derived using the skein relation. For the Kronecker type, an alternative expansion formula was given in [20], and we show that the two expansion formulas coincide.
@@ 1. INTRODUCTION
Cluster algebras were introduced by Fomin and Zelevinsky around 2000 in [9]. Later, Berenstein and Zelevinsky introduced quantum cluster algebras in [3]. The theory of cluster algebra is related to many other branches of mathematics, such as Lie theory, representation theory of algebras, Teichmüller theory and mathematical physics.
@@
Cluster algebra coming from a marked surface was studied in [8]. In this paper, We study the surface without punctures: Consider a compact connected oriented 2-dimensional bordered Riemann surface $S$ and a finite set of marked points $M$ lying on the boundary of $S$ with at least one marked point on each boundary component. In [8], a cluster algebra $\mathcal{A}(S,M)$ correspond to the marked surface $(S,M)$ with a triangulation $\Gamma$. The main results are there is a bijection $\gamma\mapsto x_{\gamma}$ between the internal arcs in $S$ and the cluster variables in $\mathcal{A}(S,M)$, furthermore the triangulations are bijective with clusters and flips are corresponding to mutations. Later [18] and [17] invented the notion of a complete path and gave the expression of $x_{\gamma}$ for the cluster algebra $\mathcal{A}(S,M)$ with principal coefficients. Later [14] gave a expansion formula by perfect matchings which is essentially the same. The expansion formula was then generalised to marked surface in [15]. The quantum cluster algebras arising from a marked surface are natural, and all the results have their counterparts in the quantum case. Then Huang gave the expansion formula for quantum cluster variables in [11].
@@
The Jacobian algebras associated with the quivers with potential given by the triangulations of surfaces are gentle algebras. The module categories of gentle algebras are studied in [2] and the cluster categories $C$ of quiver with potential are studied in [1]. A correspondence between (closed) curves on surfaces and objects in cluster categories was given by [4]. Specifically, the string objects in $C$ are indexed by curves and can be identified with string modules in the gentle algebra associate with a triangulation. Palu [16] defined a cluster character for any object in some 2-Calabi-Yau category. Later Qin [7] gave a quantum Laurent polynomial of any cluster variables using Serre polynomials.
@@
Then the connection between expansion formulas of arcs and cluster characters of modules should be studied as they both provide Laurent expansion formulas of cluster variables in the same cluster algebra. This question is addressed through a comparison between perfect matchings and canonical submodules in [22].
@@
The aim of this paper is to provide a module-theoretic interpretation for Huang’s expansion formula. The key point is to find a way to define the function $v_{\gamma}$ on canonical submodules. Huang defined a function $v$ on perfect matchings by an initial condition and a recursion formula. His formula depend on count the number of edges label some fixed notion. The correspondence between perfect matchings and canonical submodules in [22] could gave a natural way to define $v_{\gamma}$ by count the vertices of canonical submodules with some fixed label. Specifically, let $\gamma$ be an arc in $(S,M)$ with string $w$. Let $M(w)$ be the corresponding string module and $CS(M(w))$ be the set of canonical submodules.
@@
There exists $v_{\gamma}:CS(M(w))\rightarrow\mathbb{Z}$ be a map such that
@@
    If two submodules $N$ and $N^{\prime}$ with $N=N^{\prime}\cup{x_{j}}$ for an index $j$, then
@@
We show the map $v_{\gamma}$ coincides with the map $v$ in [11]. As a consequence, we obtain the following result.
@@
The quantum cluster variable asssociated with $M(w)$ can be expressed by
@@
The relationship between skein algebras and quantum cluster algebras was established in [13]. In this framework, the multiplication formulas admit a natural interpretation via the bijection between arcs and cluster variables. We explicitly describe the multiplication formulas by identifying the strings of modules that appear in the resulting expressions.
@@ 2. REVIEW OF CLUSTER ALGEBRA
We will work over a fixed algebraically closed field $K$. Let $q$ be a formal variable, $\widetilde{B}=(b_{ij})$ be a $m\times n$ matrix with full rank and $\Lambda$ be a skew-symmetric matrix satisfying
@@
It is proved in [3] that the mutation here can be applied on the new seed on any direction. If we collect all the seed which can obtained from a finite sequence of mutations on the initial seed, we can build a $n$-regular tree $\mathbb{T}_{n}$ where the vertices correspond the quantum seeds and edges correspond mutations in some direction. We denote $t_{0}$ the vertex of the initial seed. For any vertex $t$ in $\mathbb{T}_{n}$, the quantum seed $X(t)=(X_{1}(t),\cdots,X_{m}(t))$ can be obtained by a finite sequence of mutation in direction $k_{1},\cdots,k_{s}$ and we denote by $t=\mu_{k_{s}}\cdots\mu_{k_{1}}(t_{0})$.
@@
By definition, all $X_{i}(t)$ equal for any $n+1\leq i\leq m$ and $t\in\mathbb{T}_{n}$ and let $X_{n+1},\cdots,X_{m}$ denote the common value.
@@
    $X_{i}(t)$ is called quantum cluster variable for $1\leq i\leq n$ and $t\in\mathbb{T}_{n}$.
@@
    $X_{i}$ is called frozen variable for $n+1\leq i\leq m$.
@@ 3. GENTLE ALGEBRA FROM SURFACE
  • the endpoint of $\gamma$ are marked points in $M$.
@@
We call an arc $\gamma$ a boundaryarcboundary\ arc if it cuts out a digon(that is, $\gamma$ is homotopic to a curve on the boundary that intersects $M$ only on its endpoints). Otherwise, $\gamma$ is said to be an internalarcinternal\ arc. Each arc is considered up to homotopy in the class of such curves. A triangulationtriangulation of $(S,M)$ is a maximal collection $\Gamma$ of arcs that does not intersect in the interior of $S$. We only consider marked surfaces that admits a triangulation.
@@
    At each point of $Q$ start at most two arrows and stop at most two arrows.
@@
Given a string $w$, we denote by $M(w)$ the corresponding string module over $A$: the underlying vector space is obtained by replacing each vertex of $w$ by a copy of the field $K$, and the action of an arrow $a\in Q_{1}$ on $M(w)$ is the identity morphism if $a$ or its inverse is an letter of $w$, zero otherwise.
@@
The submodule $N$ of $M$ is a canonical embedding if the injective map $N\rightarrow M$ induced by the identity on the non-zero component of $N$. In this way, $N$ is called a canonical submodule of $M$. Denote $CS(M)$ the set of canonical submodules of $M$.
@@
It is clear that each substring induces a canonical submodule of $M(w)$. Moreover, since the decomposition is disjoint with maximal length, the supports of any substring modules is pairwise disjoint. So any subset in $S(w)$ induces a canonical submodule of $M(w)$. The following lemma shows the canonical submodule is equivalent to $S(w)$.
@@
From now on, we will identify a canonical submodule with its index set. Sometimes, we will use the index set to represent a string module (which may not necessarily be a submodule) whose support is the index set, provided there is no ambiguity.
@@ 4. CLUSTER EXPANSION FORMULA
In this section, we recall in [8] the cluster algebra coming from an unpunctured surface $(S,M)$. Given a triangulation $\Gamma$ of surface, let $B(\Gamma)$ denote the matrix associated with the quiver $Q(\Gamma)$. That is, $B(\Gamma)=(b_{ij})$ is an $n\times n$ matrix where:
@@
Snake graph is a significant tool in [15] to prove the positivity of the cluster algebra from surface and in [11] for quantum version. Here we recall the key result in their paper.
@@ 4.1. Snake graph and perfect matchings
For each $p_{j}$, we associate a tiletile $G(p_{j})$ as follows. Define $\Delta^{j}_{1}$ and $\Delta^{j}_{2}$ to be two triangles with edges labeled as in $\Delta_{j-1}$ and $\Delta_{j}$, further, the orientations of $\Delta^{j}_{1}$ and $\Delta^{j}_{2}$ both agree with those of $\Delta_{j-1}$ and $\Delta_{j}$ if $j$ is odd; the orientations of $\Delta^{j}_{1}$ and $\Delta^{j}_{2}$ both disagree with those of $\Delta_{j-1}$ and $\Delta_{j}$ if $j$ is even. We glue $\Delta^{j}_{1}$ and $\Delta^{j}_{2}$ at the edge labeled $\tau_{i_{j}}$, so that the orientation of $\Delta^{j}_{1}$ and $\Delta^{j}_{2}$ both either agree or disagree with those of $\Delta_{j-1}$ and $\Delta_{j}$. We say the edge labeled $\tau_{i_{j}}$ the diagonal of $G(p_{j})$.
The two arcs $\tau_{i_{j}}$ and $\tau_{i_{j+1}}$ form two edges of the triangle $\Delta$. Denote the third edge of $\Delta$ by $\tau_{[\gamma_{j}]}$. After gluing the tiles $G(p_{j})$ and $G(p_{j+1})$ at the edge labeled $\tau_{[\gamma_{j}]}$ for $1\leq j\leq d-1$ step by step, we obtain a graph, denote as $\overline{G_{\Gamma,\gamma}}$. Let $G_{\Gamma,\gamma}$ be the graph obtained from $\overline{G_{\Gamma,\gamma}}$ by removing the diagonal of each tile.
@@
    We call a snake graph $G$ with tiles $G_{1},\cdots,G_{s}$ a zigzagzigzag if for all $i$ either $G_{i}$ is glued on top of $G_{i-1}$ and $G_{i+1}$ is glued to the right of $G_{i}$ or if $G_{i}$ is glued to the right of $G_{i-1}$ and $G_{i+1}$ is glued on top of $G_{i}$
@@
    We call a snake graph $G$ with tiles $G_{1},\cdots,G_{s}$ a straightpiecestraight\ piece if for all $i$ either $G_{i}$ is glued to the right of $G_{i-1}$ and $G_{i+1}$ is glued to the right of $G_{i}$ or if $G_{i}$ is on top of $G_{i-1}$ and $G_{i+1}$ is glued on top of $G_{i}$
@@
Let $G_{i}$ and $G_{i+1}$ be two consecutive tiles of $G_{\Gamma,\gamma}$ sharing same edge a. If $b$ and $c$ are edges of $G_{i}$ and $G_{i+1}$, respectively, which are incident to $a$, then $b$ and $c$ can not simultaneously be in a perfect matching of $G$.
@@
Let $a_{1}$ and $a_{2}$ be the two edges of $G_{\Gamma,\gamma}$ which lie in counterclockwise direction from the diagonal of $G(p_{1})$. Then the minimalminimal matchingmatching $P_{-}(G_{\Gamma,\gamma})$ is defined as the unique matching which contains only boundary edges and does not contain edges $a_{1}$ or $a_{2}$. The maximalmaximal matchingmatching $P_{+}(G_{\Gamma,\gamma})$ is the other matching with only boundary edges.
@@
A perfect matching $P$ can twisttwist on a tile $G(p)$ if $G(p)$ has two edges belong to $P$. In such case we define the twisttwist $\mu_{p}(P)$ of $P$ on $G(p)$ to be the perfect matching obtained from $P$ by replacing the edges in $G(p)$ by the remaining two edges.

### v2

@@ Abstract.
We provide a module-theoretic interpretation of the expansion formula given by Huang (2022), which defines a map on perfect matchings to compute the expansion of quantum cluster variables in quantum cluster algebras arising from unpunctured surfaces. In addition, we present a multiplication formula for string modules with one-dimensional extension space, derived using the skein relations. For the Kronecker type, an alternative expansion formula was given in Çanakçı and Lampe (2020), and we show that the two expansion formulas coincide.
Keywords: Quantum cluster algebras; string modules; perfect matchings; canonical submodules; multiplication formulas.
MSC (2020): 13F60; 16G20.
@@ 1. INTRODUCTION
Cluster algebras were introduced by Fomin and Zelevinsky around 2000 in [15]. Later, Berenstein and Zelevinsky introduced quantum cluster algebras in [4]. The theory of cluster algebras is related to many other branches of mathematics, such as Lie theory, representation theory of algebras, Teichmüller theory and mathematical physics.
@@
Cluster algebras coming from marked surfaces were studied in [17]. In this paper, we study surfaces without punctures. Consider a compact connected oriented 2-dimensional bordered Riemann surface $S$ and a finite set of marked points $M$ lying on the boundary of $S$ with at least one marked point on each boundary component. In [17], a cluster algebra $\mathcal{A}(S,M)$ corresponds to the marked surface $(S,M)$ with a triangulation $\Gamma$. The main result is that there is a bijection $\gamma\mapsto x_{\gamma}$ between the internal arcs in $S$ and the cluster variables in $\mathcal{A}(S,M)$, furthermore triangulations are in bijection with clusters and flips correspond to mutations. Later, [25] and [26] invented the notion of a complete path and gave an expression of $x_{\gamma}$ for the cluster algebra $\mathcal{A}(S,M)$ with principal coefficients. Later [20] gave an expansion formula using perfect matchings which is essentially the same. The expansion formula was then generalized to marked surfaces in [21]. The quantum cluster algebras arising from a marked surface are natural, and all the results have their counterparts in the quantum case. Then Huang gave the expansion formula for quantum cluster variables in [18].
@@
The Jacobian algebras associated with the quivers with potential given by the triangulations of surfaces are gentle algebras. The module categories of gentle algebras are studied in [2] and the cluster category $C$ of a quiver with potential is studied in [1]. A correspondence between (closed) curves on surfaces and objects in cluster categories was given by [5]. Specifically, the string objects in $C$ are indexed by curves and can be identified with string modules in the gentle algebra associated with a triangulation. Palu [23] defines a cluster character for any object in some 2-Calabi-Yau category. Later, Qin [24] gives quantum Laurent polynomials for all cluster variables using Serre polynomials.
@@
Then the connection between expansion formulas for arcs and cluster characters of modules should be studied as they both provide Laurent expansion formulas of cluster variables in the same cluster algebra. This question is addressed through a comparison between perfect matchings and canonical submodules in [11]. In the quantum case, Huang introduced an expansion formula for arcs, which is defined recursively on the lattice of perfect matchings of a snake graph. The connection between perfect matchings and canonical submodules still holds. The problem is how to interpret both the recursions themselves and the values arising from them in terms of canonical submodules.
@@
The aim of this paper is to provide a module-theoretic interpretation of Huang’s expansion formula. The key point is to find a way to define the function $v_{\gamma}$ on canonical submodules. Huang defined a function $v$ on perfect matchings by an initial condition and a recursive formula. His formula depends on counting the number of edges with some fixed label. The correspondence between perfect matchings and canonical submodules in [11] could give a natural way to define $v_{\gamma}$ by counting the vertices of canonical submodules with some fixed label. Specifically, let $\gamma$ be an arc in $(S,M)$ with string $w$. Let $M(w)$ be the corresponding string module and $CS(M(w))$ be the set of canonical submodules.
@@
There exists a map $v_{\gamma}:CS(M(w))\rightarrow\mathbb{Z}$ such that
@@
    If two submodules $N$ and $N^{\prime}$ satisfy $N=N^{\prime}\cup{x_{j}}$ for some index $j$, then
@@
We show that the map $v_{\gamma}$ coincides with the map $v$ in [18]. As a consequence, we obtain the following result.
@@
The quantum cluster variable associated with $M(w)$ can be expressed by
@@
Cluster multiplication formulas play an important role in cluster theory and have attracted significant attention. In the context of acyclic cluster algebras, Sherman and Zelevinsky [27] first established such formulas for rank 2 cluster algebras of finite and affine types, which were later extended to the affine type $A_{2}^{(1)}$ by Cerulli [13]. Caldero and Keller [8] then constructed cluster multiplication formulas for finite types. This result was subsequently generalized to acyclic types by Xiao and Xu [28], as well as Xu [29]. In the framework of acyclic quantum cluster algebras, Ding and Xu [14] first obtained cluster multiplication formulas for the Kronecker case. Later, Bai, Chen, Ding, and Xu [3] extended this result to the affine type $A_{2}^{(2)}$.
The relationship between skein algebras and quantum cluster algebras was established in [22]. In this framework, the multiplication formulas admit a natural interpretation via the bijection between arcs and cluster variables. We explicitly describe the multiplication formulas by identifying the strings of the modules appearing in the resulting expressions.
@@ 2. REVIEW OF CLUSTER ALGEBRA
We will work over a fixed algebraically closed field $K$. Let $q$ be a formal variable, $\widetilde{B}=(b_{ij})$ be a $m\times n$ matrix with full rank, and $\Lambda$ be a skew-symmetric matrix satisfying
@@
It is proved in [4] that the mutation here can be applied on the new seed on any direction. If we collect all the seeds which can be obtained from a finite sequence of mutations on the initial seed, we can build an $n$-regular tree $\mathbb{T}_{n}$ where the vertices correspond to the quantum seeds and edges correspond to mutations in some direction. We denote $t_{0}$ the vertex of the initial seed. For any vertex $t$ in $\mathbb{T}_{n}$, the quantum seed $X(t)=(X_{1}(t),\cdots,X_{m}(t))$ can be obtained by a finite sequence of mutations in direction $k_{1},\cdots,k_{s}$ and we denote by $t=\mu_{k_{s}}\cdots\mu_{k_{1}}(t_{0})$.
@@
By definition, all $X_{i}(t)$ are equal for any $n+1\leq i\leq m$ and $t\in\mathbb{T}_{n}$ and let $X_{n+1},\cdots,X_{m}$ denote the common value.
@@
    $X_{i}(t)$ is called a quantum cluster variable for $1\leq i\leq n$ and $t\in\mathbb{T}_{n}$.
@@
    $X_{i}$ is called a frozen variable for $n+1\leq i\leq m$.
@@ 3. GENTLE ALGEBRA FROM SURFACE
  • the endpoints of $\gamma$ are marked points in $M$.
@@
We call an arc $\gamma$ a boundaryarcboundary\ arc if it cuts out a digon (that is, $\gamma$ is homotopic to a curve on the boundary that intersects $M$ only on its endpoints). Otherwise, $\gamma$ is said to be an internalarcinternal\ arc. Each arc is considered up to homotopy in the class of such curves. A triangulationtriangulation of $(S,M)$ is a maximal collection $\Gamma$ of arcs that does not intersect in the interior of $S$. We only consider marked surfaces that admit a triangulation.
@@
    At each vertex of $Q$, at most two arrows start and at most two arrows stop.
@@
Given a string $w$, we denote by $M(w)$ the corresponding string module over $A$: the underlying vector space is obtained by replacing each vertex of $w$ by a copy of the field $K$, and the action of an arrow $a\in Q_{1}$ on $M(w)$ is the identity morphism if $a$ or its inverse is a letter of $w$, and zero otherwise.
@@
The submodule $N$ of $M$ is a canonical embedding if the injective map $N\rightarrow M$ is induced by the identity on the non-zero component of $N$. In this way, $N$ is called a canonical submodule of $M$. Denote by $CS(M)$ the set of canonical submodules of $M$.
@@
It is clear that each substring induces a canonical submodule of $M(w)$. Moreover, since the decomposition is disjoint with maximal length, the supports of the substring modules are pairwise disjoint. So any subset of $S(w)$ induces a canonical submodule of $M(w)$. The following lemma shows that canonical submodules correspond to elements of $S(w)$.
@@
From now on, we will identify a canonical submodule with its index set. Sometimes, we will use the index set to represent a string module (which need not necessarily be a submodule) whose support is the index set, provided there is no ambiguity.
@@ 4. CLUSTER EXPANSION FORMULA
In this section, we recall from [17] the cluster algebra coming from an unpunctured surface $(S,M)$. Given a triangulation $\Gamma$ of a surface, let $B(\Gamma)$ denote the matrix associated with the quiver $Q(\Gamma)$. That is, $B(\Gamma)=(b_{ij})$ is an $n\times n$ matrix where:
@@
The snake graph is a significant tool in [21] to prove positivity for cluster algebras from surfaces, and in [18] for the quantum version. Here we recall the key result in their paper.
@@ 4.1. Snake graph and perfect matchings
For each $p_{j}$, we associate a tiletile $G(p_{j})$ as follows. Define $\Delta^{j}_{1}$ and $\Delta^{j}_{2}$ to be two triangles with edges labeled as in $\Delta_{j-1}$ and $\Delta_{j}$, further, the orientations of $\Delta^{j}_{1}$ and $\Delta^{j}_{2}$ both agree with those of $\Delta_{j-1}$ and $\Delta_{j}$ if $j$ is odd; the orientations of $\Delta^{j}_{1}$ and $\Delta^{j}_{2}$ both disagree with those of $\Delta_{j-1}$ and $\Delta_{j}$ if $j$ is even. We glue $\Delta^{j}_{1}$ and $\Delta^{j}_{2}$ at the edge labeled $\tau_{i_{j}}$, so that the orientation of $\Delta^{j}_{1}$ and $\Delta^{j}_{2}$ both either agree or disagree with those of $\Delta_{j-1}$ and $\Delta_{j}$. We call the edge labeled $\tau_{i_{j}}$ the diagonal of $G(p_{j})$.
The two arcs $\tau_{i_{j}}$ and $\tau_{i_{j+1}}$ form two edges of the triangle $\Delta$. Denote the third edge of $\Delta$ by $\tau_{[\gamma_{j}]}$. After gluing the tiles $G(p_{j})$ and $G(p_{j+1})$ at the edge labeled $\tau_{[\gamma_{j}]}$ for $1\leq j\leq d-1$ step by step, we obtain a graph, denoted by $\overline{G_{\Gamma,\gamma}}$. Let $G_{\Gamma,\gamma}$ be the graph obtained from $\overline{G_{\Gamma,\gamma}}$ by removing the diagonal of each tile.
@@
    We call a snake graph $G$ with tiles $G_{1},\cdots,G_{s}$ a zigzagzigzag if for all $i$ either $G_{i}$ is glued on top of $G_{i-1}$ and $G_{i+1}$ is glued to the right of $G_{i}$ or if $G_{i}$ is glued to the right of $G_{i-1}$ and $G_{i+1}$ is glued on top of $G_{i}$.
@@
    We call a snake graph $G$ with tiles $G_{1},\cdots,G_{s}$ a straightpiecestraight\ piece if for all $i$ either $G_{i}$ is glued to the right of $G_{i-1}$ and $G_{i+1}$ is glued to the right of $G_{i}$ or if $G_{i}$ is on top of $G_{i-1}$ and $G_{i+1}$ is glued on top of $G_{i}$.
@@
Let $G_{i}$ and $G_{i+1}$ be two consecutive tiles of $G_{\Gamma,\gamma}$ sharing the same edge a. If $b$ and $c$ are edges of $G_{i}$ and $G_{i+1}$, respectively, which are incident to $a$, then $b$ and $c$ can not simultaneously be in a perfect matching of $G$.
@@
Let $a_{1}$ and $a_{2}$ be the two edges of $G_{\Gamma,\gamma}$ which lie in the counterclockwise direction from the diagonal of $G(p_{1})$. Then the minimalminimal matchingmatching $P_{-}(G_{\Gamma,\gamma})$ is defined as the unique matching which contains only boundary edges and does not contain edges $a_{1}$ or $a_{2}$. The maximalmaximal matchingmatching $P_{+}(G_{\Gamma,\gamma})$ is the other matching with only boundary edges.
@@
A perfect matching $P$ can twisttwist on a tile $G(p)$ if $G(p)$ has two edges belonging to $P$. In such a case we define the twisttwist $\mu_{p}(P)$ of $P$ on $G(p)$ to be the perfect matching obtained from $P$ by replacing the edges in $G(p)$ by the remaining two edges.